OpenStudy (anonymous):

Find the x- and y-intercept of the line. x + 4y = 36? help

OpenStudy (jdoe0001):

notice this line  there are 2 points there on point hits the y-axis, thus called the y-intercept another point hits the x-axis, thus called the x-intercept notice the one hitting the "y", y = 2, what is "x"? notice the one hitting the "x", x = 4, what is "y"?

OpenStudy (anonymous):

so what do i do after that?

OpenStudy (anonymous):

x-intercept is 9; y-intercept is 36. is that the answer?

OpenStudy (jdoe0001):

well, notice when the line hits the "y", the x = 0 when it hits the "x", y =0 how to find the intercepts? if you want the x-intercept, set \(\bf y = 0\) in the equation and solve for "x" if you want the y-intercept, set \(\bf x = 0\) in the equation and solve for "y"

OpenStudy (anonymous):

x+4y=36 set y=0 first to get your x-int x+4(0)=36 x=36 then set x=0 to get y-int 0+4y=36 y=9

OpenStudy (anonymous):

thanks jojo, i juss had to switch them right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Write y = mc012-1.jpgx + 4 in standard form using integers.\[Write y \frac{ 3 }{ 5 }x + 4 \in standard form using integers.? help jojo!\]

OpenStudy (anonymous):

add y= to the second equation

OpenStudy (anonymous):

well standard form all it means that your equation should look in this form ex. x+y=0 meaning all the variables on one side so for this one move 3/5x to the side where y is... ill show you y=35x+4∈ y-35x=4∈

OpenStudy (anonymous):

–3x – 5y = 20 would that be the answer?